RC Update and Run plan for Sep 18,2020

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are M-F at 4:15 PM

To join the Meeting:  https://bluejeans.com/861439920
To join via phone :  +1.408.740.7256 (US (San Jose)), +1.888.240.2560 (US Toll Free))

Friday 18 Sept 2020.

General updates

(See d2n Run Plan 2020 for broader status/overview.)

  • Lost RF at 11:00am on Thursday [1], down until 4pm
    • Root cause was issue with the Master Oscillator (MO) that drives timing for all RF cavities and other Accelerator systems.. [2]
  • Continued production data in Kin-C, Kin-Z with pol 3He target, Transverse/270 when beam returned.
  • 10:30pm set HMS to Kin-A (HMS Kin-A : p0=-4.200, Theta=13.50) and continued production running there
    • Fairly smooth beam overnight
  • 6am Began the transition to 1-pass
    • RadCon enetered at 6am and we were in CA by 6:45
    • Reconfigured the Raster for 1-pass [3], [4], [5]
    • Reset the SHMS (and HMS) collimator controls and rehomed/reset both to correct collimators [6]
      • Everything behaving normally after the controller reset.
    • Opened He3 cylinder valves at target [7]
    • Install collimator in target; buttoned up target [8]
    • Change spectrometer angles to SHMS: 8.5, HMS:11.7 and set momenta appropriately
    • Hall A setup the charge-asym feedback system for us (P. King, et al)
  • Hall to Beam Permit at 10:30am; MCC started sending tune beam around 11:15
    • Some issues with beam focus (too narrow in y), got things squared away by 1pm
  • Laser interlock tripped around 1pm (coincidentally)
    • Took access and found a panel had shifted (tape relaxing) [9]; Fixed and back to beam permit around 2pm.
  • Initial target polarization after target work ~25% [10]
    • Updated NMR calibration constant estimate: 7.39 [11], [12], [13]
  • Beam energy measurement started at 2:15pm [14]
    • Results: 2182.708 +/- 2.3662 MeV [15], [16]
  • Will continue with 1-pass runplan

Notes for the Future

  • Charge Feedback Parity DAQ Instructions will need to be followed by Shift Crew -- We will have to run the 'flipper script'
  • 1-pass running program
  • Hall will move to Controlled Access on Monday
    • When is still TBD; depends when RadCon is available
    • Could be in the afternoon
    • EPR photodiode will be replaced on Monday (JP, Junhao, RadCon)
    • Will move to Restricted Access ________ ?
  • Bill, Murchhana check compass/field measurements
    • Also checking with Jixie re: modeled fields
    • Need to make a write-up that clearly defines all the different coordinate systems used, swapped power supply/polarity issues, etc.

Pending Issues

Opportunistic Jobs/Access

Current Run Plan

Thursday DAY → Thursday SWING

  • Continue production data in Kin-C, Kin-Z with pol 3He target, Transverse/270 polarization with pNMR/NMR every 5 hrs.
    • All runs for Kin-C and Kin-Z settings use the T1 trigger only
  • Set HMS only to Kin-A mid-SWING (HMS Kin-A : p0=-4.200, Theta=13.50)
    • Run T1 only on SHMS (as usual)
    • Run T1 (1 hour), T2 (3 hours) on HMS; then repeat that cycle until morning

Friday OWL

  • Continue production data in Kin-A, Kin-Z with pol 3He target, Transverse/270 polarization with pNMR/NMR every 5 hrs.
    • Run T1 only on SHMS (as usual)
    • Run T1 (1 hour), T2 (3 hours) on HMS; and repeat that cycle until morning
  • 6 am Friday Morning
    • Keys will be tight -- no 'extra' people in the Hall please
    • Ensure RadCon is ready to enter Hall
    • RadCon will survey Hall and begin transition to Controlled Access
    • Stop beam, take NMR/pNMR
    • Rotate target to Longitudinal/180, switch on longitudinal lasers, set QWP and Correction Coils appropriately
      • Continue to pump in longitudinal direction until 6:30pm or so (ie. RadCon is done Survey and Hall is in regular Controlled Access
    • Shut off lasers and lock-out fibers
      • All personnel that will be on/near the pivot (RadCon, Techs, and Target personnel) must attach their LOTO locks -- Please remind those going downstairs

Friday DAY

  • (Started at 6am) Open target enclosure
    • Install collimators (JP, RadCon)
    • Check on EPR photo-diode status, likely replace it? (JP, RadCon)
  • Ensure fast raster is reconfigured correctly (Gunning, Brad)
  • Rotate to minimum angles for HMS (11.7) and SHMS (8.5) (S. Lassiter, A. Comer)
  • Reset SHMS collimator controls (Brad)
  • Ensure Qasym feedback system is visible and ready for Hall C control (Bill)
  • Switch over any software updates needed for 1-pass (Chao, Sylvester)
  • Target Configured for Longitudinal/180, QWP and correction coils set, Long. lasers are back ON (after target enclosure buttoned up)
  • Close up and wait for MCC to establish (1-pass) beam to the Hall
  • NOTE: Hall C is in control of the Charge Asymmetry feedback system until Monday
    • Charge Feedback Parity DAQ Instructions
    • We will have to run the ./FlipPCvaluesHallC script in the Parity DAQ VNC session after every IHWP change

Friday SWING → Monday

  • See task list at Run Plan for 1-pass for more details
    • Priority will be PbPt polarization cross check with longitudinal pol. elastics.
    • These have lower priority but will likely be pursued first while target spins up
      • Beam energy measurement (1st thing)
      • 3He, N2 density measurements
      • H2 xsec measurement
      • Pressure curves
  • NOTE: Hall C is in control of the Charge Asymmetry feedback system until Monday
    • Charge Feedback Parity DAQ Instructions
    • We will have to run the ./FlipPCvaluesHallC script in the Parity DAQ VNC session after every IHWP change


Standing Orders

  • Change the IHWP at least every 8 hours when running on glass cells. Call Halls A and B in advance. If there are any issues, call the RC
    • NOTE: Hall C is in control of the Charge Asymmetry feedback system until Monday Charge Feedback Parity DAQ
    • We will have to run the ./FlipPCvaluesHallC script in the Parity DAQ VNC session after every IHWP change
  • Ensure that all information in the CODA run-start GUIs is correct for every run!
  • If beam goes away for more than 3h, or MCC does operations on the beam that may change the beam tune, then move target to No Target position,
    • verify Harps scans at 3H07A and 3H07B are within spec, and
    • ensure beam is stable (with all locks ON) at 0, 0mm before continuing.
    • Beam recovery procedure has more detailed procedure if needed.
  • New beam positions have been verified and established to be (0,0) mm for all targets.
    • Monitor that beam is at (0,0) using the stripchart on the Wall screen and verify no scraping on the glass cells using the Analysis monitor.
  • Check the position locks are on.
  • Check that CODA is running normally
  • Check online replays shortly after starting runs, & full replay results too.
  • Fill BTA with realistic numbers. You may follow the guidance coming from "load from EPICS", if it does not overestimate the ABU (as it generally does). https://logbooks.jlab.org/entry/3783171. If we have to lower the beam current we actually want because the beam is too unstable, prorate the ABUs.
!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!
!!                                                     !!
!!  ATTN: SHIFT LEADERS                                !!
!!     Please follow the shift summary template on     !!
!!     the laminated SL instructions.                  !!
!!                                                     !!
!!  ATTN: TARGET OPERATORS                             !!
!!  All NMR measurement HCLOG entries should include:  !!
!!     1) A title that clearly lists polarization      !!
!!     2) Screenshots of both NMR_FieldSweep and       !!
!!        the polarization fit                         !!
!!                                                     !!
!!     For an example see the following entry:         !!
!!     https://logbooks.jlab.org/entry/3768696       !!
!!                                                     !!
!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

Notices in Effect for Shift Crew

  • Policy for CANS-related hall status change: Make a logbook entry when the hall goes in, or comes out of laser-controlled access.
  • Magnet policies
    • Follow the updated guidelines in the How-to#Spectrometer_Magnets Spectrometer Magnets How to
    • Spectrometer GUI (go_magnets) should always be visible on Shift Leader computer, hcdesk3
    • If any magnet trips, snapshot the associated power supply window and post it to hclog before resetting the interlock.
  • Monitor the beam position on target and call MCC immediately if you see large excursions (1 mm level).
  • When rotating the spectrometer, use HCCAM01 to keep an eye on the SOLID Cherenkov test stand when rotating SHMS (in addition to watching the live angle-cam image, etc). The updated rotation instructions are here: https://hallcweb.jlab.org/wiki/index.php/Spectrometer_Rotation#NOTICES_IN_EFFECT

Useful Links

PolHe3 Polarization Measurements

Misc Links